The current in a conductor and the potential difference across its ends are measured by an ammeter and a voltmeter. The meters draw negligible currents. The ammeter is accurate but the voltmeter has a zero error (that is, it does not read zero when no potential difference is applied). Calculate the zero error if the readings for two different conditions are 1.75 A, 14.4 V and 2.75 A, 22.4 V.

Grade:10

1 Answers

Kevin Nash
9 years ago
Sol. Let the voltmeter reading when, the voltage is 0 be X. l base 1R/ l base 2R = V base 1/V base 2 ⇒ 1.75/2.75 = 14.4 – V/22.4 – V ⇒ 0.35/0.55 = 14.4 – V/22.4 – V ⇒ 0.07/0.11 = 14.4 – V/22.4 – V ⇒ 7/11 = 14.4 – V/22.4 – V ⇒ 7(22.4 - V) = 11(14.4 - V) ⇒ 156.8 – 7V = 158.4 – 11V ⇒(7 - 11)V = 156.8 – 158.4 ⇒ -4V = - 1.6 ⇒ V = 0.4 V.